# Privacy-Preserving Logistic Regression Training with A Faster Gradient Variant

### v8

@@ 5 Experiments
For a fair comparison with the baseline (Kim et al., 2018a), we utilized the same 10-fold cross-validation (CV) technique on the same iDASH dataset consisting of 1579 samples with 18 features and the same 5-fold CV technique on the other five datasets. Like (Kim et al., 2018a), We consider the average accuracy and the Area Under the Curve (AUC) as the main indicators. Tables 1 and 2 display the results of the two experiments, respectively. The two tables also provide the average evaluation running time for each iteration and the storage (encrypted dataset for the baseline work and encrypted dataset and $\bar{B}$ for our method). We adopt the same packing method that Kim et al. Kim et al. (2018a) proposed and hence our solution has similar storage of ciphertexts to (Kim et al., 2018a) with some extra ciphertexts to encrypt the $\bar{B}$. We chose $1+0.9^{t}$ as our learning rate configuration.
@@ 6 Conclusion
In this paper, we proposed a faster gradient variant called quadratic gradient, and implemented the quadratic-gradient version of NAG in the encrypted domain to train the logistic regression model.
The quadratic gradient presented in this work can be constructed from the Hessian matrix directly, and thus somehow combines the first-order Newton’s method and the second-order gradient (descent) method together. There is a promising chance that quadratic gradient could accelerate other gradient methods.
Also, quadratic gradient might substitute and supersede the line-search method, for example when using enhanced Adagrad-like methods, and could use gradient descent methods to accelerate Newton’s method, resulting in super-quadratic algorithms.
It is highly likely that the quadratic gradient approach could combine first-order gradient descent/ascent algorithms with the second-order Newtonian method, and that it could be applicable to a wide range of numerical optimization problems.

### v9

@@ 5 Experiments
For a fair comparison with the baseline (Kim et al., 2018a), we utilized the same 10-fold cross-validation (CV) technique on the same iDASH dataset consisting of 1579 samples with 18 features and the same 5-fold CV technique on the other five datasets. Like (Kim et al., 2018a), We consider the average accuracy and the Area Under the Curve (AUC) as the main indicators. Tables 1 and 2 display the results of the two experiments, respectively. The two tables also provide the average evaluation running time for each iteration and the storage (encrypted dataset for the baseline work and encrypted dataset and $\bar{B}$ for our method). We adopt the same packing method that Kim et al. (2018a) proposed and hence our solution has similar storage of ciphertexts to (Kim et al., 2018a) with some extra ciphertexts to encrypt the $\bar{B}$. We chose $1+0.9^{t}$ as our learning rate configuration.
@@ 6 Conclusion
In this paper, we proposed a faster gradient variant, termed quadratic gradient, and implemented the quadratic-gradient version of NAG in the encrypted domain to train a logistic regression model. The quadratic gradient introduced in this work can be directly constructed from the Hessian, effectively merging first-order gradient (descent) methods with the second-order Newton’s method. There is a promising potential for the quadratic gradient to accelerate other gradient methods, and it may serve as an alternative or even a replacement for the traditional line-search method.